# Fenglin Liu1, Jinge Wu211footnotemark: 1, Hongjian Zhou1, Xiao Gu1, Soheila Molaei1,&Anshul Thakur1, Lei Clifton3, Honghan Wu2,4, David A. Clifton1,5

### v1

@@ Abstract
The application of Large Language Models (LLMs) to various clinical applications has attracted growing research attention. LLMs currently achieve competitive results compared to human experts in examinations. However, real-world clinical decision-making differs significantly from the standardized, exam-style scenarios commonly used in current efforts. It therefore remains a challenge to apply LLMs to complex medical tasks that require a deep understanding of medical knowledge. A common approach is to fine-tune LLMs for target tasks, which, however, not only requires substantial data and computational resources but is also still prone to generating ‘hallucinations’. In this paper, we present the RiskAgent system to perform a broad range of medical risk predictions, covering over 387 risk scenarios across diverse complex diseases, e.g., cardiovascular disease and cancer. RiskAgent is designed to collaborate with hundreds of clinical decision tools, i.e., risk calculators and scoring systems that are supported by evidence-based medicine. To evaluate our method, we have built the first benchmark MedRisk specialized for risk prediction, including 12,352 questions spanning 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems. The results show that our RiskAgent, with 8 billion model parameters, achieves 76.33% accuracy, outperforming the most recent commercial LLMs, o1, o3-mini, and GPT-4.5, and doubling the 38.39% accuracy of GPT-4o. On rare diseases, e.g., Idiopathic Pulmonary Fibrosis (IPF), RiskAgent outperforms o1 and GPT-4.5 by 27.27% and 45.46% accuracy, respectively. Finally, we further conduct a generalization evaluation on an external evidence-based diagnosis benchmark and show that our RiskAgent achieves the best results. These encouraging results demonstrate the great potential of our solution for diverse diagnosis domains. For example, instead of extensively fine-tuning LLMs for different medical tasks, our method, which collaborates with and utilizes existing evidence-based medical tools, not only achieves trustworthy results but also reduces resource costs, thus making LLMs accessible to resource-limited clinical applications. To improve the adaptability of our model in different scenarios, we have built and open-sourced a family of models ranging from 1 billion to 70 billion parameters. Our code, data, and models are all available at https://github.com/AI-in-Health/RiskAgent.
Introduction
@@
Large Language Models (LLMs), such as the GPT series [1, 2, 3, 4] and PaLM [5], have shown promising performance in understanding text and assisting humans. Inspired by the success of LLMs, an increasing amount of research is attempting to apply LLMs to the medical field and explore their potential in assisting healthcare professionals, resulting in different types of medical LLMs [6]. Specifically, based on the LLMs PaLM [5] and Gemini [7], Google has developed MedPaLM-2 [8] and Med-Gemini [9, 10], respectively, which achieve 86.5% and 91.1% accuracy comparable to the 87.0% accuracy of human experts [11] on the US Medical Licensing Examination (USMLE). Based on open-source LLM LLaMA [12, 13], dozens of medical LLMs have been proposed, such as Clinical Camel [14] and Meditron [15], to address different medical tasks.
@@
Figure 1: a. The RiskAgent, including three LLM agents (Decider, Executor, and Reviewer) can perform multiple medical risk predictions given the patient’s healthcare information. b. Statistics of the MedRisk benchmark, which consists of 12,352 medical risk questions, covering 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems. c. With only 8 billion parameters, our RiskAgent outperforms both existing high-performance general LLMs (i.e., GPT-4o [3], o1 [4], o3-mini [16]) and state-of-the-art medical LLM (i.e., Meditron-70B) by large margins across different diseases, symptoms, specialties, and organ systems. In contrast to existing LLMs, RiskAgent can collaborate with evidence-based medical tools to not only substantially increase its risk prediction accuracy, but also deliver evidence-based answers. The t-tests between the results from RiskAgent and the best-performing LLMs indicate that the improvement is significant with $p<0.01$. d. The examples of risk predictions by our method for cancer, cardiac events, and asthma, demonstrate greater accuracy than GPT-4o. The pink- and blue-colored text indicates the incorrect and desirable answers, respectively.
@@
However, the effective application of LLMs to complex medical tasks remains challenging: 1) Clinical inefficacy: LLMs achieve superior performance mainly in examination tasks [17]. However, existing LLMs, including GPT-4, perform poorly on tasks closer to clinical decision-making, such as medical code querying [18] and new drug understanding [17]. 2) Resource-intensive: To enable LLMs to deliver accurate results on different tasks, a potential solution is fine-tuning LLMs on target medical data [19, 6]. However, fine-tuning LLMs (forcing them to learn all tasks) not only places significant learning pressure on them, resulting in decreased performance, but also inevitably requires substantial data (ranging from 1 million [20, 21, 8] to 80 billion tokens [15, 11]) and computational resources that may be prohibitively difficult to obtain. More importantly, these fine-tuned medical LLMs do not utilize or collaborate with various tools already deployed in hospitals, including devices, models, and APIs, leading to resource waste. 3) Privacy concern: An alternative solution is to prompt commercial LLMs (e.g., the GPT series [22]). However, the use of commercial LLMs involves strict privacy regulations regarding sensitive patient information in the real world, raising privacy concerns for their adoption in hospital scenarios [23]. 4) Unfaithful outputs: LLMs still face the widely known ‘hallucination’ challenge [24, 25, 26]: the produced answers appear reasonable but are not based on factual information and knowledge. Furthermore, existing LLMs could not effectively and accurately provide evidence to show the sources of generated answers, which is crucial in healthcare [27, 6]. Therefore, developing different medical LLMs for diverse medical tasks is expensive, time-consuming, and energy-intensive.
@@
In this work, we present RiskAgent for solving complex medical problems and take risk prediction as a representative example that requires LLMs to not only accurately understand complex patients’ health records, but also predict potential health risks, including the risk of developing diseases or the mortality/survival rate of diseases. This is critical for preventative health [28, 29], since early and accurate risk prediction can alert physicians and patients for early intervention, and thus improve clinical outcomes (e.g., increasing patient survival), especially for complex diseases such as cardiovascular disease and rare diseases [30, 31]. As shown in Figure 1(a), our RiskAgent includes three LLM agents: Decider, Executor, and Reviewer: 1) The Decider analyzes medical problems and accurately selects appropriate tools from hundreds of options; 2) The Executor understands the selected tools, parses their required parameters, and executes them; 3) The Decider then analyzes the returned tool outputs and provides initial answers; 4) The Executor structures and executes outputs to generate the final answers; 5) The Reviewer finally reviews the decision-making process and provides reflection on the results. To comprehensively evaluate LLMs’ risk prediction performance for diverse scenarios, we build a risk prediction benchmark MedRisk including 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems, totaling 12,352 cases. Figure 1(b) presents the benchmark’s statistics. We evaluate 13 state-of-the-art methods, covering both general and medical LLMs, as well as open-source public and commercial LLMs. Figure 1(c) and Table 1 reveal that current LLMs perform poorly in making accurate risk predictions (15.83$\sim$58.77accuracy) for complex diseases, while our RiskAgent substantially outperforms existing LLMs with large margins. Figure 1(d) shows that our method makes more accurate risk predictions than GPT-4o [3] for different complex diseases. The first two GPT-4o examples show that incorrectly predicting high-risk patients as low-risk would lead to delayed treatment. Conversely, when low-risk patients are predicted to be high-risk, as shown in the third GPT-4o example, it may cause over-treatment, waste medical resources, and bring unnecessary anxiety to patients.
Overall, our RiskAgent can make a wide range of risk predictions for hundreds of scenarios across diverse diseases. Our RiskAgent has the following advantages: 1) Clinical efficacy: The objective of an LLM is to learn to collaborate with existing medical tools supported by evidence-based medicine [32], rather than learning to perform various complex medical tasks. This greatly reduces LLM’s learning pressure and makes it easier to deliver superior performance and, therefore, is more efficient and straightforward than learning from a large volume of target task data as in conventional methods. 2) Resource efficiency: In real-world settings, particularly hospitals, there are existing deployed medical tools, including devices, models, and APIs. Our method learns to collaborate with and directly utilize these tools, instead of designing and training new medical LLMs for clinical tasks, thus reducing resource costs and making our LLMs accessible for resource-limited clinical applications. This is a critical step toward sustainable AI. 3) Privacy friendly: The encouraging results of our model, with only 8 billion parameters, show the potential of our solution to build open-source LLMs in the real world with fewer privacy concerns compared to commercial LLMs. 4) Faithful outputs: Compared to existing methods, the collaboration with evidence-based medical tools not only brings more accurate performance, but also, most importantly, helps our model produce trustworthy results with means for clinicians to efficiently inspect the evidence of the output recommendations, which is critical in achieving trust from clinical users.
Framework
@@
Figure 2 shows the flowchart of our RiskAgent system. Inspired by the success of actor-critic reinforcement learning [33] that introduces the Critic, Actor, Reward, and Environment, we introduce three LLM agents and one component, i.e., Decider, Executor, Reviewer, and Environment. These modules work together to ensure that our system runs smoothly and accurately.
@@
  • Executor: The LLM is instructed to execute the Decider’s decisions, i.e., parse the required parameters to invoke the tools, and convert the answers into risk scores or choices for evaluation.
@@
  • Environment: It provides the system with external information that can be interacted with, such as the medical tool library, required parameters for the tools, tool outputs, and system prompts.
@@
We further introduce the Reviewer that evaluates the provided answer by considering the historical analysis process and providing reflection results to improve the system’s performance. If the Reviewer determines that the answer is correct, i.e., Reflect[right], the system outputs the answer; otherwise, i.e., Reflect[wrong], the Reviewers provides reflection results to the Reviewer and Executor, and the system continues running. In the case of the Decider selecting the wrong tools, the Reviewer provides reflection results on the Decider’s decisions, and the Decider re-selects the tools. The Reviewer also handles the situation when the Executor fails to correctly implement the Decider’s decisions, such as a parameter parsing failure. The Reviewer, therefore, also reviews the progress of the Executor’s work to ensure the tool invocation is completed successfully.
Environment
We introduce five environments to enhance collaboration between our system and medical tools, shown in Figure 2 (b).
  • Environment 1 Hundreds of medical tools are often involved in real-world clinical scenarios, due to their complexity and diversity. To enable LLMs to collaborate with these tools, we introduce a retrieval-ranking algorithm to extract the $N$ most relevant medical tools from a tool library containing $M$ tools based on the current medical problem.
  • Environment 2 When the Decider has chosen a tool, this environment provides the Executor with the required parameters for tool invocation, enabling the Executor to use the tools smoothly and accurately.
  • Environment 3 The returned results from the medical tools are stored in this environment, where the Decider interacts with for analysis and the generation of free-text answers to questions.
  • Environment 4 This environment provides the system prompts to assist the Executor in formatting and converting the free-text answers generated by the Decider into risk scores or choices for evaluation.
  • Environment 5 This environment stores the historical analysis process from the Decider and Executor, as well as system prompts that instruct the Reviewer to perform the review process accurately.
We attach a detailed description of our method in our supplementary material (Appendix A).
Results
In this section, we comprehensively evaluate the effectiveness of our solution and compare it with current representative LLMs, covering both state-of-the-art closed-source commercial LLMs and open-source public LLMs with model parameters ranging from 7 to 70 billion.
Types Methods MedRisk-Qualitative
@@ Results
Public LLMs Mistral-7B [34] 47.89 29.75 33.72 36.29 35.47
LLaMA-3-8B [35] 48.59 44.62 42.59 46.33 43.75
Mixtral-8x7B [36] 44.37 29.43 34.19 37.07 34.42
LLaMA-3-70B [35] 52.82 48.42 48.31 53.28 48.62
PMC-LLaMA-13B [11] 29.58 18.04 17.27 21.24 18.83
Meditron-70B [15] 35.21 24.68 26.02 28.96 27.19
Commer. LLMs GPT-4o [3] 58.17(1.76) 52.28(1.01) 50.15(0.16) 53.28(0.98) 50.68(0.20)
o3-mini [4] 58.59(2.41) 53.61(0.79) 52.16(0.60) 52.59(1.02) 52.55(0.60)
o1 [4] 61.27 56.33 56.59 55.60 56.25
GPT-4.5 [37] 57.04 55.38 53.79 57.53 54.22
AI Agents ReAct-8B [38] 52.82 53.80 50.18 49.81 51.14
BOLAA-8B [39] 55.63 56.65 55.43 52.90 52.60
Chameleon-8B [40] 57.04 57.91 53.56 55.21 54.30
RiskAgent-8B 74.09(2.06) 76.84(1.05) 78.11(0.72) 71.51(0.75) 78.34(0.43)
Types Methods MedRisk-Quantitative
@@
Public LLMs Mistral-7B [34] 21.13 23.73 25.55 22.39 24.35
LLaMA-3-8B [35] 28.87 27.85 32.79 32.05 32.06
Mixtral-8x7B [36] 16.90 17.09 19.84 17.37 18.83
LLaMA-3-70B [35] 32.39 31.96 35.59 38.22 34.42
PMC-LLaMA-13B [11] 21.13 13.92 15.29 16.99 15.83
Meditron-70B [15] 26.76 24.05 24.50 24.32 24.59
Commer. LLMs GPT-4o [3] 31.69(3.80) 35.51(0.88) 39.49(0.49) 39.54(2.02) 38.39(0.79)
o3-mini [4] 51.97(3.34) 59.11(0.99) 55.54(0.97) 61.70(4.64) 55.91(1.02)
o1 [4] 60.56 58.23 58.46 61.00 58.77
GPT-4.5 [37] 39.44 37.66 39.79 42.47 39.04
AI Agents ReAct-8B [38] 34.51 40.82 44.57 35.91 40.50
BOLAA-8B [39] 40.85 47.47 50.41 46.33 46.02
Chameleon-8B [40] 38.73 50.32 48.42 49.81 47.89
RiskAgent-8B 77.74(1.58) 76.01(0.31) 75.61(0.86) 70.50(0.93) 76.33(0.73)
Table 1: Performances of different methods on risk predictions. ‘B’: Billion. We use accuracy to report our results. We conduct five runs for our RiskAgent, GPT-4o, and o3-mini. We report the mean and standard deviation${}_{\text{(STD)}}$ of performance. In this paper, all values are reported in percentage (%). Higher is better for all columns. The bold number denotes the best result across all methods. We select the most common cases across disease, symptom, specialty, and organ system for a broad evaluation and subsequently report the overall results (i.e., the average performance on all test data).
Figure 3: Performance for RiskAgent-8B, GPT-4o, and o3-mini. In the boxplot, the central line indicates the median value, while the lower and upper boundaries represent the 25th (Q1) and 75th (Q3) percentiles, respectively. The whiskers extend up to 1.5 times the interquartile range (IQR).

### v2

# RiskAgent: Synergizing Language Models with Validated Tools for Evidence-Based Risk Prediction

@@ Abstract
Large Language Models (LLMs) achieve competitive results compared to human experts in medical examinations. However, it remains a challenge to apply LLMs to complex clinical decision-making, which requires a deep understanding of medical knowledge and differs from the standardized, exam-style scenarios commonly used in current efforts. A common approach is to fine-tune LLMs for target tasks, which, however, not only requires substantial data and computational resources but also remains prone to generating ‘hallucinations’. In this work, we present RiskAgent, which synergizes language models with hundreds of validated clinical decision tools supported by evidence-based medicine, to provide generalizable and faithful recommendations. Our experiments show that RiskAgent not only achieves superior performance on a broad range of clinical risk predictions across diverse scenarios and diseases, but also demonstrates robust generalization in tool learning on the external MedCalc-Bench dataset, as well as in medical reasoning and question answering on three representative benchmarks—MedQA, MedMCQA, and MMLU.
Machine Learning, ICML
1 Introduction
@@
Inspired by the success of Large Language Models (LLMs) (Hurst et al., 2024; Jaech et al., 2024; Chowdhery et al., 2022), an increasing amount of research is attempting to apply LLMs to the medical field and explore their potential in assisting healthcare professionals, resulting in different types of medical LLMs (Zhou et al., 2023). For example, based on the LLMs PaLM (Chowdhery et al., 2022) and Gemini (Google, 2023), Google has developed MedPaLM-2 (Singhal et al., 2023) and Med-Gemini (Yang et al., 2024), respectively, which achieve 86.5% and 91.1% accuracy comparable to the 87.0% accuracy of human experts (Wu et al., 2024) on the US Medical Licensing Examination. Based on open-source LLM LLaMA (Touvron et al., 2023a, b), dozens of medical LLMs have been proposed, such as Clinical Camel (Toma et al., 2023) and Meditron (Chen et al., 2023), to address different medical tasks.
@@
Figure 1: a. The RiskAgent, including three LLM agents (Decider, Executor, and Reviewer) can perform multiple medical risk predictions given the patient’s healthcare information. b. Statistics of the MedRisk benchmark, which consists of 12,352 medical risk questions, covering 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems. c. With only 8 billion parameters, our RiskAgent outperforms both existing high-performance general LLMs (i.e., GPT-4o (Hurst et al., 2024), o1 (Jaech et al., 2024), o3-mini (gpt, 2025a)) and state-of-the-art medical LLM (i.e., Meditron-70B) by large margins across different diseases, symptoms, specialties, and organ systems. In contrast to existing LLMs, RiskAgent can collaborate with evidence-based medical tools to not only substantially increase its risk prediction accuracy, but also deliver evidence-based answers. The t-tests between the results from RiskAgent and the best-performing LLMs indicate that the improvement is significant with $p<0.01$. d. The examples of risk predictions by our method for cancer, cardiac events, and asthma, demonstrate greater accuracy than GPT-4o. The pink- and blue-colored text indicates the incorrect and desirable answers, respectively. The correct answers are verified by clinicians.
@@
However, the effective application of LLMs to complex medical tasks remains challenging: 1) Clinical inefficacy: LLMs achieve superior performance mainly in examination tasks (Liu et al., 2024). However, existing LLMs, including GPT-4, perform poorly on tasks closer to clinical decision-making, such as medical code querying (Soroush et al., 2024); 2) Resource-intensive: To enable LLMs to deliver accurate results on different tasks, a solution is fine-tuning LLMs on target medical data (Van Veen et al., 2024; Zhou et al., 2023). However, fine-tuning LLMs (forcing them to learn all tasks) not only places significant learning pressure on them, but also inevitably requires substantial data (ranging from 1 million (Ferber et al., 2024; Han et al., 2023; Singhal et al., 2023) to 80 billion tokens (Chen et al., 2023; Wu et al., 2024)) and computational resources that may be prohibitively difficult to obtain for resource-limited medical institutions; 3) Privacy concern: An alternative solution is to prompt commercial LLMs (e.g., the GPT series (Nori et al., 2023)). However, the use of commercial LLMs involves strict privacy regulations regarding sensitive patient information in the real world (Ullah et al., 2024). 4) Unfaithful outputs: LLMs still face the widely known ‘hallucination’ challenge (Pal et al., 2023; Ji et al., 2023): the produced answers appear reasonable but are not based on factual information and knowledge. Furthermore, existing LLMs could not effectively and accurately provide evidence to show the sources of generated answers, which is crucial in healthcare (Kitamura, 2023; Zhou et al., 2023). Therefore, developing different medical LLMs for diverse medical tasks is expensive, time-consuming, and energy-intensive.
@@
In this work, we present RiskAgent for solving complex medical problems and take risk prediction as a representative example that requires LLMs to not only accurately understand complex patients’ health records, but also predict potential health risks, including the risk of developing diseases or the mortality/survival rate of diseases. This is critical for preventative health (Pudjihartono et al., 2022), since early and accurate risk prediction can alert physicians and patients for early intervention, and thus improve clinical outcomes, especially for complex diseases such as cardiovascular disease and rare diseases (Damen et al., 2016). Figure 1(a) shows that our RiskAgent includes three LLM agents: Decider, Executor, and Reviewer: 1) The Decider analyzes medical problems and accurately selects appropriate tools from hundreds of options; 2) The Executor understands the selected tools, parses their required parameters, and executes them; 3) The Decider then analyzes the returned tool outputs and provides initial answers; 4) The Executor structures and executes outputs to generate final answers; 5) The Reviewer finally reviews the decision-making process and provides reflection on the results.
[…]
Overall, the main contributions of our work are:
  • We present RiskAgent, a clinically efficient multi-agent framework that synergizes language models with diverse validated tools to provide faithful, generalizable, and evidence-based recommendations.
  • We construct MedRisk, a comprehensive benchmark for generalist medical risk prediction. It encompasses 12,352 distinct clinical questions spanning 154 diseases, 86 symptoms, and 50 specialties.
  • We conduct extensive experiments showing that RiskAgent outperforms state-of-the-art models and exhibits generalization capabilities in risk prediction, tool learning, medical reasoning, and question answering.
2 Related Work
Medical Large Language Models
The adaptation of general LLMs to the medical domain has gained significant attention (Zhou et al., 2023). While these models demonstrate proficiency in standardized examinations, they frequently struggle with real-world clinical tasks requiring high fidelity and precision (Soroush et al., 2024). A significant limitation is their reliance on internal parametric knowledge for tasks that demand precise calculations or adherence to updated clinical guidelines. This approach ignores the utility of existing deployed medical tools—such as devices, models, and APIs—leading not only to resource waste but also to potential hallucinations and a disconnect from evidence-based practice (Sackett, 1997). Instead of attempting to embed all medical knowledge into model weights, our work focuses on enabling LLMs to collaborate with validated evidence-based medical tools.
Tool-Augmented Medical Agents
Recent works have explored augmenting medical LLMs with external tools to enhance clinical capabilities. 1) Multi-agent reasoning: MedAgents (Tang et al., 2024) and MDAgents (Kim et al., 2024) simulate clinical discussions by coordinating multiple agents with different medical roles to improve diagnostic reasoning. While effective for complex decision-making, these systems still rely on parametric knowledge without leveraging validated clinical calculators. 2) General tool calling: Some approaches equip medical LLMs with general-purpose tools such as code interpreters or search engines to assist clinical tasks. However, these generic tools lack the domain-specific validation required for precise medical risk assessment. 3) Clinical calculator integration: AgentMD (Jin et al., 2024) pioneers connecting LLMs with clinical calculators, demonstrating the value of evidence-based tools. However, it focuses on limited tool coverage and relies on extensive prompting of proprietary models. Our RiskAgent advances this direction by training a lightweight model that collaborates with hundreds of validated tools, reducing resource costs while producing faithful outputs that clinicians can efficiently verify.
3 Approach
@@
Figure 2 shows the flowchart of our RiskAgent system [^1] We attach a detailed description of our method in Appendix A.. Inspired by the success of actor-critic reinforcement learning (Grondman et al., 2012) that introduces the Critic, Actor, Reward, and Environment, we introduce three LLM agents and one component, i.e., Decider, Executor, Reviewer, and Environment. These modules work together to ensure that our system runs smoothly and accurately.
3.1 LLM Agents
@@
Executor: The LLM is instructed to execute the Decider’s decisions, i.e., parse the required parameters to invoke the tools, and convert the answers into risk scores or choices.
@@
We further introduce the Reviewer that evaluates the provided answer by considering the historical analysis process and providing reflection results to improve the system’s performance. If the Reviewer determines that the answer is correct, i.e., Reflect[right], the system outputs the answer; otherwise, i.e., Reflect[wrong], the Reviewers provides reflection results to the Reviewer and Executor, and the system continues running. In the case of the Decider selecting the wrong tools, the Reviewer provides reflection results on the Decider’s decisions, and the Decider re-selects the tools. The Reviewer also handles the situation when the Executor fails to correctly implement the Decider’s decisions, such as a parameter parsing failure. The Reviewer also reviews the progress of the Executor’s work to ensure the tool invocation is completed successfully.
3.2 Environment
It provides the system with external information that can be interacted with, such as the medical tool library, required parameters for the tools, tool outputs, and system prompts. We introduce five environments to enhance synergy between LLMs and tools. Environment 1: Hundreds of medical tools are often involved in real-world clinical scenarios, due to their complexity and diversity. To enable LLMs to collaborate with these tools, we introduce a retrieval-ranking algorithm to extract the $N$ most relevant medical tools from a tool library containing $M$ tools based on the current medical problem. Environment 2: When the Decider has chosen a tool, this environment provides the Executor with the required parameters for tool invocation, enabling the Executor to use the tools smoothly and accurately. Environment 3: The returned results from the medical tools are stored in this environment, where the Decider interacts with for analysis and the generation of free-text answers to questions. Environment 4: This environment provides the system prompts to assist the Executor in formatting and converting the free-text answers generated by the Decider into risk scores or choices for evaluation. Environment 5: This environment stores the historical analysis process from the Decider and Executor, as well as system prompts that instruct the Reviewer to perform the review process accurately.
[…]
Types Methods MedRisk-Qualitative MedRisk-Quantitative
@@ 3.2 Environment
Public LLMs Mistral-7B (Jiang et al., 2023) 47.89 29.75 33.72 36.29 35.47 21.13 23.73 25.55 22.39 24.35
LLaMA-3-8B (AI@Meta, 2024) 48.59 44.62 42.59 46.33 43.75 28.87 27.85 32.79 32.05 32.06
Mixtral-8x7B (Jiang et al., 2024) 44.37 29.43 34.19 37.07 34.42 16.90 17.09 19.84 17.37 18.83
LLaMA-3-70B (AI@Meta, 2024) 52.82 48.42 48.31 53.28 48.62 32.39 31.96 35.59 38.22 34.42
PMC-LLaMA-13B (Wu et al., 2024) 29.58 18.04 17.27 21.24 18.83 21.13 13.92 15.29 16.99 15.83
Meditron-70B (Chen et al., 2023) 35.21 24.68 26.02 28.96 27.19 26.76 24.05 24.50 24.32 24.59
Commer. LLMs GPT-4o (Hurst et al., 2024) 58.17(1.76) 52.28(1.01) 50.15(0.16) 53.28(0.98) 50.68(0.20) 31.69(3.80) 35.51(0.88) 39.49(0.49) 39.54(2.02) 38.39(0.79)
o3-mini (Jaech et al., 2024) 58.59(2.41) 53.61(0.79) 52.16(0.60) 52.59(1.02) 52.55(0.60) 51.97(3.34) 59.11(0.99) 55.54(0.97) 61.70(4.64) 55.91(1.02)
o1 (Jaech et al., 2024) 61.27 56.33 56.59 55.60 56.25 60.56 58.23 58.46 61.00 58.77
GPT-4.5 (gpt, 2025b) 57.04 55.38 53.79 57.53 54.22 39.44 37.66 39.79 42.47 39.04
AI Agents ReAct-8B (Yao et al., 2023) 52.82 53.80 50.18 49.81 51.14 34.51 40.82 44.57 35.91 40.50
BOLAA-8B (Liu et al., 2023) 55.63 56.65 55.43 52.90 52.60 40.85 47.47 50.41 46.33 46.02
Chameleon-8B (Lu et al., 2024) 57.04 57.91 53.56 55.21 54.30 38.73 50.32 48.42 49.81 47.89
RiskAgent-8B 74.09(2.06) 76.84(1.05) 78.11(0.72) 71.51(0.75) 78.34(0.43) 77.74(1.58) 76.01(0.31) 75.61(0.86) 70.50(0.93) 76.33(0.73)
3.3 Instruction Fine-tuning
Following the workflow established in Environments 1-5, we construct samples with explicit instructions and well-defined output formats. Each sample includes: (a) patient information, clinical queries, and candidate tools, with the expected output being the selection of the correct tool with justification; (b) patient information and tool schema, with the expected output being correctly formatted parameter extraction; (c) the original results computed by calculators with the sample question, with the expected output being a rephrased answer to the question; (d) outputs from step c along with questions and options, with the expected output being the final selected option; and (e) information from steps a-d, with the expected output being a comprehensive reflection (the instruction template is shown in Figure 5). We take the LLaMA-3-8B LLM (AI@Meta, 2024) as our backbone to train our model using parameter-efficient fine-tuning techniques (LoRA) (Hu et al., 2022), significantly reducing computational requirements while maintaining high performance across medical risk prediction tasks.
4 Experiments
In this section, we evaluate the effectiveness of our solution for risk prediction, tool learning, and question answering  [^2] We attach a detailed description of our settings in Appendix A.
4.1 Evaluation on Risk Prediction